OpenStudy (anonymous):

antiderivative of 4rootx-(4/rootx)

OpenStudy (anonymous):

\[4\sqrt{x}-\frac{ 4 }{ \sqrt{x} }\]

OpenStudy (anonymous):

[4((2/3x^(3/2))-(2x^(1/2)))\]